OpenStudy (anonymous):

what is the mclaurin series expansion of 1/sqrt(1-(x^2))

OpenStudy (anonymous):

guess we can grind it til we find it i cannot think of a snappy shortcut for this one

OpenStudy (anonymous):

\[f(0)=1\] so it start with 1 then \(f'(x)=\frac{x}{(1-x^2)^{\frac{3}{2}}}\) making \(c_1=0\)

OpenStudy (anonymous):

in fact the function is even, so only even exponents

OpenStudy (anonymous):

i cannot think of a snappy way to do this, sorry i guess you just have to keep taking derivatives oh wait, maybe there is a quick way!!

OpenStudy (anonymous):

\[\frac{1}{\sqrt{1-x^2}}\] is the derivative of \(\sin^{-1}(x)\) so if you know the maclaurin for that one, you can differentiate term by term to get this one